Calculate the molarity of each of the following solutions: (a) 30 g of Co(NO3)2. 6H2O in
4.3 L of solution (b) 30 mL of 0.5 M H2SO4 diluted to 500 mL.

Grade:12

3 Answers

Vikas TU
10 years ago

 

APPROVE!

Bevkoof Singh
10 years ago

1-2.015

2-

SWAMY
4 years ago
Here,
The initial molarity or the M1 is 0.5M.
And the initial volume or the V1 is 30 ml.

And,
The final volume or the V2 is 500 ml.
So the,
Final volume will be,
M1V1=M2V2
0.5×30=M2×500
M2=0.5×30/500.
M2=0.03

The final molarity is 0.03 M